A beam of parallel light, 1.08mm in diameter passes through a lens with a focal length of 11.3cm. Another lens, this one of focal length 19.8cm, is located behind the first lens so that the light traveling out from it is again parallel.

A. What is the distance between the two lenses? Answer given is 31.1cm

B. How wide is the outgoing beam? Answer given is 1.89mm
Answer #1

a)the distance between the two lenses=f1+f2

=11.3+19.8

=31.1 cm

b)magnification=f1/f2

=19.8/11.3

=1.7522

so d=1.08*m

=1.08*1.7522

=1.89 mm